### v1

Abstract
Scientific Workflow Systems (SWSs) are advanced software frameworks that drive modern research by orchestrating complex computational tasks and managing extensive data pipelines. These systems offer a range of essential features, including modularity, abstraction, customization, interoperability, workflow composition tools, resource management, error handling, and comprehensive documentation. Utilizing these frameworks accelerates the development of scientific computing, resulting in more efficient and reproducible research outcomes. Despite their significance, developing a user-friendly, efficient, and adaptable SWS poses several challenges that are not always well-documented or understood. This study explores these challenges through an in-depth analysis of interactions on Stack Overflow (SO) and GitHub, key platforms where developers and researchers discuss and resolve issues. In particular, we leverage topic modeling (BERTopic) to understand the topics SWSs developers discuss on these platforms. Then, we examine the popularity and difficulty of those topics. We identified 10 topics developers discuss on SO (e.g., Workflow Creation and Scheduling, Data Structures and Operations, Workflow Execution) and found that workflow execution is the most challenging among them. By analyzing GitHub issues, we identified 13 topics (e.g., Errors and Bug Fixing, Documentation, Dependencies) and discovered that data structures and operations is the most difficult topics in this context. We also found common topics between SO and GitHub, such as data structures and operations, task management, and workflow scheduling. Additionally, we categorized each topic by type (How, Why, What, and Others) and observed that the How type consistently dominates across all topics, indicating a need for procedural guidance among developers. This dominance of the How type is also prevalent in other domains, such as Chatbots and Mobile development. We believe that our study will guide future research in proposing tools and techniques to help the community overcome the challenges developers face when developing SWSs.
keywords:
Scientific Workflow Systems, Stack Overflow, GitHub, Topic Modeling, Developer Challenges
1 Introduction
Scientific Workflow Systems (SWSs) are software frameworks designed to streamline the design, execution, and management of complex, data-intensive scientific experiments and research [1]. They allow users to create, visualize, and automate workflows composed of interconnected steps, each carrying out specific data operations. These systems orchestrate tasks, manage data dependencies, and ensure reproducibility by consistently capturing and executing workflows [2]. By enhancing the efficiency of running and analyzing experiments, they facilitate collaboration and meet the rigorous demands of scientific research. SWSs serve as the bedrock upon which efficient, transparent, and impactful research is built, propelling knowledge and innovation across diverse fields. They accomplish mundane yet essential tasks intrinsic to scientific inquiry—ranging from data acquisition to analysis and reporting. By liberating researchers from time-consuming tasks, SWSs allow them to focus on more complex and creative work, which helps speed up and improve research results [3]. They provide a platform to create and distribute the necessary computational steps for tasks involving data analysis and simulations, all while concealing the intricate infrastructural complexities [4]. More importantly, they capture the entire scientific experimentation process, offering valuable insights for reproducing, reusing, or adapting these procedures.
Traditionally, various types of scientific analyses are implemented by practitioners using either scripting languages (Python, Perl, or Bash) requiring fairly sophisticated computational skills or by manually calling and storing intermediate results, which requires a considerable amount of time organizing data files and managing the overall analysis process [5]. Alternatively, SWSs can help practitioners by providing higher-level modeling approaches for explicitly specifying analyses as well as by providing several generic services for optimizing workflow execution, visualizing workflows and workflow results, managing intermediate data products, and storing the details of past workflow executed results [6]. Their ability to handle large volumes of data and integrate various computational tools makes them essential for advancing research across diverse scientific domains. Furthermore, SWSs afford researchers, developers, and project managers real-time insights into the progress of research endeavors, empowering them to identify bottlenecks, allocate resources judiciously, and optimize workflow execution. The customizable nature of SWSs allows practitioners to tailor workflows to their specific needs and preferences, enhancing their utility and versatility. In essence, SWSs emerge not merely as tools of convenience but as indispensable allies in the relentless pursuit of scientific excellence.
Figure 1: Year-wise Number of Posts, Issues, and Pull Requests
The extensive adoption of SWSs in diverse fields (e.g., machine learning [7], data mining [8], bioinformatics [9], chemistry [10], geosciences [11], high-performance computing [12]) has led the software engineering community to develop numerous software tools tailored for SWSs. Thousands of developers are contributing to SWSs in various capacities, undertaking tasks such as system architecture design (determining the overall structure, components, and interactions between different modules to ensure flexibility, scalability, and performance), algorithm development (developing algorithms for data processing, analysis, and automation, leveraging techniques from machine learning, natural language processing, and data mining to extract insights from scientific data), UI design and development (user-friendly interfaces, i.e., intuitive dashboards, visualization tools, and workflow editors), backend development (backend logic of SWSs, including data storage, processing pipelines, and workflow execution engines, writing codes in different languages and integrating with databases and external services), integration with external tools and services (interoperability and data exchange). They are also involved in testing and quality assurance, documentation, deployment and maintenance, security and compliance, and performance optimization by identifying bottlenecks, tuning algorithms and configurations, and leveraging parallel processing and distributed computing techniques to efficiently handle large-scale data analysis. Managing these vast amounts of tasks comes with numerous challenges. Developers usually turn to SO to ask questions, find solutions, learn and understand new technologies or programming languages, validate best practices, and engage with the community [13]. GitHub also plays a crucial role in collaborative development [14]. Through GitHub issues and pull requests, developers can report problems, suggest improvements, engage in discussions, contribute to projects, and participate in the open-source community. Such technical Q&As, issue reports, and pull requests may communicate developers’ challenges when developing SWSs.
Practitioners encountered a significant deficiency in usability, effectiveness, and reliability when dealing with these SWSs [15]. Furthermore, the consistent rise in the number of posts, issues, and pull requests (for selected SWSs) shown in Figure 1 implies a growing demand for increased support for developers engaged in SWSs. Given the widespread use of SWSs across various domains and the interest they have garnered, including from the software analytics community, it becomes crucial to comprehend these challenges for advancing the efficiency and effectiveness of these systems, which are integral to data-intensive research and complex scientific computations. Practitioners often encounter unique difficulties in SWSs, such as managing heterogeneous data sources, ensuring reproducibility, and optimizing workflows for high-performance computing environments. By systematically identifying these challenges, researchers and practitioners can tailor tools and methodologies to better support the development of robust, scalable, and user-friendly workflows. Addressing these challenges not only enhances the usability and reliability of SWSs but also accelerates scientific discovery by enabling researchers to focus on their domain-specific tasks rather than the intricacies of the workflow system itself. Moreover, understanding these challenges provides valuable insights that can drive innovation in SWSs design, leading to more efficient computational solutions that meet the evolving needs of the scientific community [7]. Therefore, we have analyzed SO posts, GitHub issues, and pull requests related to SWSs to identify developer challenges in our research.
In this paper, we conduct an empirical investigation to rigorously understand the needs and desiderata of developers working with SWSs. Our objective is to gain a precise understanding of the challenges developers face and identify potential areas for improvement, aiming to pinpoint specific issues and obstacles encountered in their work. To achieve this, we have compiled a dataset of 35,619 SO posts and 140,400 GitHub issues and pull requests related to SWSs. Using a combination of manual analysis and automated topic modeling, we examine these SO posts and GitHub data to uncover the specific challenges developers encounter. In particular, our study aims to answer the following three research questions (RQs):
  • RQ1: What topics are SWSs developers asking about on Stack Overflow, and which topics are raised in the GitHub projects? The SWSs-related posts, issues, and pull requests may reflect developers’ challenges when learning or developing an SWS. To understand these challenges, we utilize topic modeling to extract semantic topics from SWSs-related data on SO and GitHub. From SO posts, we identified 10 topics, which include traditional software engineering topics (e.g., data structures and operations, dependencies management) and SWSs-specific themes (e.g., workflow creation and scheduling, workflow execution). Similarly, analysis of GitHub data yielded 13 topics, encompassing traditional software engineering topics (e.g., errors and bug fixing, dependencies) and SWSs-specific topics (e.g., data management and task scheduling, annotations). By comparing the challenges developers face on SO and GitHub, we aim to identify any similarities. By answering RQ1, we seek to pinpoint the areas in SWSs development where developers frequently seek assistance, thereby highlighting the aspects that commonly pose challenges for them.
  • RQ2: What types of questions are SWSs developers asking? Developers ask different (i.e., How, Why) types of questions to address different challenges. To answer RQ2, we adopt a methodology similar to previous research on identifying types of posts on Stack Overflow [16, 17, 18]. We select a statistically significant number of posts for each topic and classify them into types such as How, What, Why, and Others. This approach enables us to systematically analyze the nature of the questions developers ask and gain deeper insights into their needs and challenges. Our analysis shows that SWSs developers primarily seek guidance on specific implementation routines, workflow examples, and troubleshooting—highlighting the dominance of How questions, similar to other domains like ChatBot and Mobile Development. Our findings underscore the need for step-by-step instructions and real-world scenarios to help developers achieve their goals in SWSs implementation.
  • RQ3: To what extent do developers perceive the revealed challenges in terms of difficulties? To assess the difficulty of each topic, we use two metrics: the percentage of unresolved posts/issues/pull requests and the median time to resolve them. This approach follows the methodologies of prior studies [16, 19, 20, 21, 22, 17]. In RQ3, our goal is to identify the most challenging topics to resolve, as these are likely to pose significant difficulties for all but the most experienced developers. We identify workflow execution as the most challenging topic on SO due to its complexity in managing distributed resources, handling large-scale data processing, ensuring fault tolerance, and optimizing performance across diverse environments. Similarly, on GitHub, Kubernetes deployments for SWSs frameworks pose significant challenges, primarily due to difficulties in ensuring scalability, fault tolerance, and efficient resource allocation required for managing complex and distributed workflows. This analysis highlights areas within SWSs development where additional support and resources are essential.
Paper Organization: The remainder of this paper is structured as outlined as follows: In Section 2, we briefly discuss Scientific Workflows, Scientific Workflow Systems(SWSs), and related work. Moving on to Section 3, we detail our study’s methodology. Our findings are presented in Section 4, and Section 5 illustrates the implications. Section 6 addresses potential threats to the validity of our results. Finally, Section 7 concludes the paper, highlighting directions for future research.
2 Background and Related Work
SWSs emerged as a response to the growing complexity and scale of scientific research [23]. As computational power increased and data generation exploded across various scientific domains, researchers needed more efficient ways to manage, automate, and reproduce their experiments. Traditional manual methods were inadequate for handling the intricate sequences of data processing tasks required. SWSs provide a structured framework for designing, executing, and sharing multi-step computational processes. They integrate diverse tools and data sources, facilitating collaboration and enhancing the reproducibility of scientific research. We aim to understand developers’ challenges in developing SWSs by analyzing SO posts, GitHub issues, and pull requests. In this section, we provide an overview of the background and previous research associated with SWSs development. Additionally, we explore prior work that utilized topic modeling techniques to gain insights into developer perspectives.
2.1 Scientific Workflow Systems
An SWS is a specialized software platform designed to automate, manage, and execute complex sequences of computational or data processing tasks, often referred to as workflows, in scientific research whose execution order is driven by a computerized representation of the workflow logic [24]. Researchers categorized existing SWSs into five distinct niches, each representing a specific approach to handling workflows in scientific computing [25]. Below, we describe these niches and provide highlighted examples of systems that exemplify each category.
First, SWSs like KNIME [26], Galaxy [27], Apache Taverna [9], and Watchdog [28] offer graphical user interfaces for the composition and execution of workflows. The shallow learning curve makes these systems accessible to everyone, even those without programming skills. Second, with systems like Anduril [29], Balsam [30], SciPipe [31], and JUDI [32], workflows are defined using a set of classes and functions in generic programming languages such as Python, Scala, and others. These systems have the advantage of being usable without a graphical interface (e.g., in a server environment) and allow workflows to be easily managed with version control systems like Git (https://git-scm.com). Third, SWSs like Snakemake [33], Nextflow [34], Bpipe [35], ClusterFlow [36], and BigDataScript [37] use domain-specific language (DSL) to specify workflows. These DSLs offer the advantages of the prior category while enhancing readability by modeling key workflow management components directly, reducing the need for extraneous operators or boilerplate code. For Nextflow and Snakemake, which extend generic programming languages (Groovy and Python), the full power of the underlying language is retained, allowing for features like conditional execution and configuration handling. Fourth, systems like Popper [38] use purely declarative approaches for workflow specification through configuration file formats like YAML [39]. While offering the concision and clarity of the third niche, these systems are exceptionally readable for non-developers. Fifth, system-independent workflow specification languages like CWL [40] and WDL [41] use declarative syntax for defining workflows, which can be executed by various executors (e.g., Cromwell (https://cromwell.readthedocs.io/en/latest/), and Toil [42]). While these languages, like those in the fourth niche, often lack integration with imperative or functional programming, they offer the advantage of executing the same workflow across different execution backends, enhancing scalability. Additionally, they facilitate interoperability among various workflow languages.
SWSs encompass several key aspects: workflow design, specification, execution, monitoring and control, resource management, data management, reproducibility and sharing, and collaboration. These systems significantly enhance productivity, collaboration, and reliability in scientific research while optimizing the efficient use of computational resources.
2.2 Prior Studies on SWSs
Numerous studies have been conducted to improve the usability, reliability, and scalability of SWSs. For example, Molder et al. [25] focus on sustainable data analysis with Snakemake. Reproducibility challenges are explored in [43], while [1] discourse reusability issues. Collaborative enhancements are discussed in [44], and Jagla et al. [45] extend SWSs for next-generation sequencing analysis. Evaluation criteria for SWSs are examined in [46, 47]. Additionally, Di Tommaso et al. address reliability in Nextflow [34], and studies on common workflow language (CWL) are discussed in [40, 48]. Jain et al. discuss SWSs for high-throughput applications in [49], while Ahmad et al. explore workflow scheduling challenges in [50]. Hundreds of SWSs have been developed to aid scientific workflow operations [51], and extensive research ([52, 53, 54, 55, 15]) has been conducted to identify the challenges associated with the efficient construction and utilization of SWSs. Several studies [56, 57, 58] have explored the intersection of SWSs and large language models (LLMs). Integrating LLMs, such as GPT models, into SWSs can drive substantial advancements across various domains, particularly in automating workflow design, optimizing data-driven processes, and enhancing scientific collaboration. This cutting-edge AI and SWSs combination could lead to more efficient, accessible, and scalable workflows. Unlike previous studies, this work makes the first attempt to understand the challenges of SWSs as perceived by developers.
Figure 2: An example of a scientific workflow using Galaxy SWS
2.3 Scientific Workflows
Modern scientific collaborations have opened up the opportunity to solve complex problems that demand multidisciplinary expertise and large-scale computational experimentation. These experiments typically involve a sequence of processing steps that need to be executed on selected computing platforms. However, executing these experiments can be difficult due to the complexity and variety of applications, the diversity of analysis goals, the heterogeneity of computing platforms, and the volume and distribution of data. A common strategy to make these experiments more manageable is to model them as workflows and to use an SWS to organize their execution [59]. Lin et al. [24] defined scientific workflow as a computerized facilitation or automation of a scientific process, in whole or part, which usually streamlines a collection of scientific tasks with data channels and dataflow constructs to automate data computation and analysis to enable and accelerate scientific discovery.
Figure 3: An example of a scientific workflow using Snakemake SWS
Practitioners can share their scientific workflows (from next, we consider workflow as scientific workflows) in several workflow repositories (e.g., WorkflowHub [60], myExperiment [61], Snakemake workflows repository [62]). These repositories contain thousands of workflows using many SWSs. Figure 2 presents a simple graphic-based workflow that explores the Iris dataset using Galaxy SWS. This workflow performs statistical analysis and creates scatterplots to visualize relationships between its features. By following these steps, we can gain a comprehensive understanding of the Iris dataset’s structure and the relationships between its features.
We show a script-based (DSL) workflow using Snakemake SWS that explores the Iris dataset from Kaggle in Figure 3. This workflow downloads the dataset, performs basic exploratory data analysis (EDA), and generates a summary report.
2.4 Topic Modeling
Topic modeling is a statistical modeling technique used to uncover the underlying structure and main themes within a large collection of documents, making it easier to organize, search, and understand the content [63]. Its applications span various text analysis tasks [64], including document clustering [65], information retrieval [66], recommendation systems [67], language translation [68], and content summarization [69]. Researchers have extensively explored topic modeling, publishing numerous articles in software engineering, political science, medicine, and linguistics. Extensive utilization of topic modeling is observed in the software engineering research community [70], particularly in domains like mining software repositories [71, 72, 73, 74, 75, 76, 77, 78], source code analysis [79, 80, 81, 82, 83, 78], spam detection [84], and recommendation systems [85, 86, 87, 88, 89], among others. While numerous papers contribute to this field, noting all of them is beyond our scope. Hence, we highlight a selection of significant papers. Several methods, such as BERTopic [90], Latent Dirichlet Allocation (LDA) [91], Non-Negative Matrix Factorization (NMF) [92], Latent Semantic Analysis (LSA) [93], Correlated Topic Model (CTM) [94], Dynamic Topic Models [95], Word Embedding-Based Models [95], and Biterm Topic Model (BTM) [96], contribute to the advancements of topic modeling.
Many of the previous studies have employed the LDA algorithm [91] or its variants for topic extraction. However, recent research has increasingly turned to BERTopic [90] for similar analyses, such as the analysis of scientific papers [97], spam detection [84], research trends on language models [98, 99], natural language processing [100], and the integration of machine learning [101]. While LDA has been a foundational and widely used method in topic modeling, it comes with significant limitations. For instance, LDA requires the number of topics to be fixed in advance, which can be difficult to determine and often necessitates manual tuning. Its reliance on a bag-of-words representation overlooks word order and context, leading to the potential loss of crucial semantic information. Moreover, as the size of the dataset or the number of topics increases, LDA’s computational demands rise, making it less scalable. The algorithm is also prone to generating incoherent topics, especially in the presence of noisy or extensive datasets, and its performance is highly sensitive to hyperparameter settings, requiring extensive experimentation.
In contrast, BERTopic [90] addresses these limitations with several advanced features. It automatically determines the number of topics, captures contextual information through transformer-based models like BERT, and is capable of handling larger datasets. BERTopic also employs hierarchical clustering and dynamic topic modeling, allowing it to adapt to changes in the data and generate more granular and interpretable topics. These strengths make BERTopic particularly well-suited for applications where a deep semantic understanding is essential. Given the short and highly technical nature of SO posts, GitHub issues, and pull requests, BERTopic is the best choice for these data. Its ability to capture nuanced semantics makes it superior for identifying contextually relevant topics. Therefore, this study plans to utilize BERTopic [90] to leverage these strengths.
2.5 Topic Analysis of Technical Q&As
Topic modeling has significant applications in technical Q&A data, facilitating the analysis and categorization of large volumes of questions and answers. By identifying underlying themes and topics within the content, topic modeling helps to organize and index data, making it easier for users to find relevant information. For instance, in forums such as Stack Overflow, topic modeling can cluster similar questions, aiding in quicker retrieval of solutions and reducing redundancy [102]. Additionally, it can track trending topics, monitor user interests, and identify common issues or knowledge gaps, proving valuable for both users seeking information and developers aiming to improve their products or services [103]. Topic models are extensively used in many studies to understand the topics of general SO posts and track topic trends [104, 105, 106, 107]. Prior work also leverages topic modeling in specific application development domains, such as mobile application development [16], machine learning application development [108], concurrency development [109], security-related development [20], DevOps development [110], chatbot development [17], classification of database technology [111], and understanding nonfunctional requirements [112]. To the best of our knowledge, topic analysis of SWSs-related posts on SO is still pending, but SWS is a promising field of software engineering. Therefore, we recognize the need to identify the topics of SWSs, and we utilize BERTopic [90] because of its contextual understanding capacity and high accuracy.
2.6 Topic Analysis of GitHub Data
Topic modeling with GitHub data involves analyzing text from repositories, issues, pull requests, and comments to identify prevalent themes or topics. This technique can uncover trends, common issues, or popular features within software development communities. Researchers apply algorithms like LDA, BERTopic, and others to categorize the vast amount of unstructured text data on GitHub, helping developers and organizations better understand user needs, project focus, and collaboration patterns. Previous studies have utilized topic modeling to analyze GitHub data, revealing prominent challenges developers face, user requirements, and emerging trends in software projects. This insight assists project managers in task prioritization and resource allocation while promoting streamlined collaboration and well-informed decision-making. By delving into the dynamics of software development communities, researchers contribute to improving software engineering methodologies. For example, Scoccia et al. [22] used LDA topic modeling to identify challenges faced by web developers, correlating real-time GitHub issues with queries on SO to explore common areas in web app development. Similarly, Li et al. [21] applied LDA topic modeling to understand challenges in quantum software engineering by analyzing SO posts and GitHub issue reports. Several other studies [113, 114, 115, 75, 116, 117] have utilized topic modeling with GitHub datasets. Given that GitHub issues and pull requests are often brief and context-specific, BERTopic’s embedding-based approach is particularly well-suited for analyzing this data. Therefore, we choose to utilize BERTopic [90] for our analysis.
3 Study Design
Figure 4: Overview of the methodology of our study
The main goal of our study is to examine what SWSs developers are asking about. To achieve this goal, we analyze discussions on SO posts, and issues and pull requests on GitHub. These platforms provide rich datasets, frequently used in research across various domains, including quantum software engineering [21], cryptography APIs [118], concurrency [109], and deep learning [119]. SO offers structured data with questions, answers, and metadata (e.g., accepted answers, scores) but lacks fine-grained topic information related to SWSs. Hence, we first need to identify posts from SO related to SWSs, group them according to their dominant topic, and then conduct our analysis. For GitHub, we focus on projects related to SWSs identified from our SO data, ensuring consistency across both platforms in the projects we analyze. Figure 4 shows the overall study design of our research. Below, we describe the study design approaches of our work.
3.1 Stack Overflow Data Collection and Prepossessing
Stack Overflow is a widely used technical Q&A forum among developers. Therefore, we chose it as one of our data sources. Below, we outline the steps involved in collecting and preprocessing SO data.
Step 1: Gather Scientific Workflow Systems List. Researchers typically use tag-based search functionality on SO to retrieve data, as authors tag their posts with tags (e.g., javascript, python, arrays) to enhance visibility and increase the possibility of receiving answers [120]. However, the field of SWSs spans multiple disciplines (e.g., bioinformatics [121], geospatial analysis [122], software engineering [123], machine learning [124], grid computing [125]), leading to diverse terminologies used by different users. This terminological inconsistency presents a challenge when attempting to conduct searches related to SWSs on SO. Additionally, SWSs have different names (e.g., Galaxy [27], NextFlow [34], SnakeMake [33], Dask [126]) depending on the specific domain or community, further complicating the process of locating these systems through specific tags on SO. Thus, we plan to gather the SWSs list and use it to identify the posts related to them. We compiled a list of SWSs by referencing multiple sources [127, 128, 129, 130]. Upon merging these lists, we identified overlaps, necessitating the removal of duplicates to ensure the integrity and completeness of our compilation. We also added the web link and GitHub link of each SWS wherever possible. Finally, we listed 352 SWS, which can be accessed using the replication package [51].
Step 2: Collecting SWSs-related posts from Stack Overflow. We collected posts from SO using the Stack Exchange Data Explorer [131] for the period between August 2008 and January 2024. As there are no predefined tags available for searching SWSs-related posts, we searched using each system’s name, gathered in the Step 1. We extracted data by using SWS names within Tags, Title, and Body. Consider, for instance, an SWS, celery. Initially, we filtered posts using the Tags celery and subsequently scrutinized whether any titles or bodies contained the term celery (this involved utilizing an SQL query to extract the dataset, such as SELECT * FROM posts WHERE LOWER(Title) LIKE ’%celery%’). Throughout this process, we exclusively focused on questions (identified by posttypeid=1). We identified posts tagged with celery that also featured celery in their Title or Body. In such instances, we selected only one post to eliminate redundancy in our dataset. In this way, we obtained 9628 posts about celery SWS by January 31, 2024.
We took precautions against potentially noisy data to ensure the accuracy of our dataset and remove any irrelevant information. Despite the possibility of encountering posts with the term celery in their title or body unrelated to the celery workflow system, we conducted a meticulous filtering process. By initially narrowing down our dataset using celery tags, we identified 9,499 posts specifically discussing the celery workflow system. This validation was reinforced by cross-referencing the tags’ descriptions provided by SO. Subsequently, we were left with 139 posts, each potentially containing content unrelated to the celery SWS. These posts varied, with some mentioning the term celery in contexts unrelated to our focus, such as referring to the vegetable, a project name, a part of a filename, or a list ID. To meticulously ensure the precision of our dataset, each of these remaining 139 posts underwent manual scrutiny. Initially, one of the authors, having over four years of experience with SWSs, meticulously reviewed each post, categorizing them as either related or unrelated to the celery workflow system. Following this initial review, two authors with over eight years of experience further scrutinized the posts. Only upon reaching a consensus did we include any posts in our final dataset. This rigorous process enabled us to identify and exclude six posts unrelated to the celery workflow system (provided in the replication package). Consequently, we finalized our dataset with 9622 posts exclusively pertinent to the celery workflow system.
In some cases, we encountered posts where all content was directly pertinent to SWSs. For instance, in the case of SnakeMake, our extraction process yielded a total of 1,696 posts, meticulously considering tags, Title, and Body. Remarkably, 1,695 of these posts were tagged with snakemake SWS, solidifying their relevance to the SnakeMake SWS. The one remaining post not tagged with SnakeMake also discussed the SnakeMake SWS, further affirming its alignment with our focus. Similar outcomes were observed for KNIME. A thorough investigation led us to identify a grand total of 308 posts related to KNIME, with 305 of them bearing the knime tags. The remaining three, lacking the knime tags, were inherently related to KNIME, underscoring their inclusion in our dataset.
Utilizing similar approaches, we explored the Galaxy SWS related posts. Our search encompassed Tags, Title, and Body, yielding a substantial total of 5652 posts. However, upon closer examination, we noted many posts tagged with android that were unrelated to Galaxy SWS. Consequently, we opted to discard these posts, amounting to the exclusion of 4378 entries. Subsequently, from the remaining 1274 posts, we further refined our dataset by filtering out posts tagged with samsung, mobile, or containing these terms in their title and body. This process led to the removal of 510 additional posts. With 764 posts remaining, we then focused specifically on those related to Galaxy SWS by filtering with the term ansible, associated with Galaxy. This step resulted in the identification of 165 posts relevant to Galaxy. Continuing our scrutiny, we observed 26 posts concerning PHP, 35 posts discussing Java, and 77 posts referencing Galaxy S4-S7. These posts were deemed irrelevant to Galaxy SWS and were consequently excluded. From the remaining 461 posts, one author meticulously assessed each post, determining their relevance to Galaxy SWS. Subsequently, the other two authors verified these assessments. Only posts that were unanimously agreed upon as relevant to Galaxy SWS were retained. Ultimately, this rigorous vetting process yielded ten posts directly related to Galaxy SWS. Thus, our finalized dataset comprises a total of 175 posts aligned with Galaxy SWS.
Table 1: Posts and Issues & Pull Requests Distribution of SWSs
SL. SWS Name Total Posts Filtered Posts No. of Issues& Pull Requests
1 Airflow 10602 10589 33163
2 SnakeMake 1696 1696 6187
3 Celery 9628 9622 13302
4 NextFlow 419 418 5397
5 KNIME 308 308 153
6 Galaxy 5648 175 39887
7 Oozie 2066 2052 674
8 Drake 846 846 20791
9 Luigi 362 348 3269
10 Dask 4681 4662 28611
11 NiFi 5187 5187 8440
For airflow, we obtained 10602 posts and we found 10448 posts with airflow tags and 31 posts having apache in title and body. Then, from the remaining 123 posts we manually checked each posts and found 13 posts unrelated with Airflow SWS. They are mainly talking about heat transfer problems, airflow temperature, respiratory airflow, function name etc. Thus, finally we have 10589 posts related with Airflow SWS.
As there are hundreds of SWS, describing each SWS is beyond our scope, though our approach is identical to that described above. In Table 1, we show the SWSs, total posts obtained after the query, and total filtered posts. For simplicity, we display the SWSs where we found more than 100 posts for a particular SWS. Our further analysis is also based on the dataset shown in Table 1. These systems are publicly available state-of-the-art systems designed for a broad spectrum of scientific analysis tasks. We gather a total of 41,443 posts. After filtration, we identified 5,540 posts unrelated to the scientific workflow system, and the remaining 35,903 posts are about SWS.
We identified several posts discussing multiple SWSs, causing the same post to appear across various workflow platforms. As illustrated in Figure 5, one such instance involves a post discussing three SWS: Airflow, Dask, and Celery.
Figure 5: A post discussed about three SWSs
As our data-gathering process queries each SWS individually, these posts are visible across all three systems. To maintain impartiality in our analysis, we selected only one instance of such posts for our examination, and we finally gathered 35,619 unique posts related to SWS. The collected data are available in the replication package [51].
Step 3: Preprocessing Scientific Workflow Systems’ posts. We first filtered out the irrelevant information before applying topic modeling techniques. While the post Title concisely summarizes the question, the Body provides crucial context and details that help us accurately identify the topic being discussed. Therefore, in this analysis, we focused on both the post Title and Body. However, the Body content might introduce noise to our analysis, so we removed quotes, HTML tags, links, and any code snippets using regular expressions from the Body of the posts. We also remove stopwords, commonly used words in the English language, such as ’how’, ’can’, and ’at’ that do not significantly affect the meaning of a sentence and can introduce bias. We used the NLTK stopwords list [132] for this operation. Then, we performed lemmatization, which involves reducing a word to its canonical form called a lemma, taking into account the word’s linguistic context (for example, the word ’scale’ is the lemma of the word ’scaling’). All the data preprocessing steps and processed data are provided in the replication package [51].
3.2 GitHub Data Collection and Prepossessing
We utilized the same SWSs listed in Table 1 for our analysis on GitHub. To gather data on issues and pull requests, we employed the GitHub REST API [133]. We observed that several SWSs are distributed across multiple repositories during this process. For instance, the Galaxy workflow system spans several repositories (e.g., tools-iuc, planemo, galaxy-hub) related to its ecosystem. Similarly, Nextflow is associated with multiple repositories (e.g., patterns, rnaseq-nf, nf-validation). Consequently, we aggregated all issues and pull requests from these relevant repositories. We treated each issue report and pull request as an individual document in the topic model. We joined the title and body of each issue and pull requests as a single document. We remove code snippets, URLs, and images from the body. Additionally, we eliminate stop words, non-alphabetical characters, and punctuation to refine the textual data. As of January 31, 2024, after preprocessing, our dataset comprises 140,400 issues and pull requests collected from various GitHub repositories. In our replication package [51], we provided the GitHub link to the corresponding SWS, along with the data collection scripts and the entire dataset.
3.3 Identify Scientific Workflow Systems’ topics.
BERTopic starts with transforming input documents into numerical representations. While there are several methods to accomplish this, SentenceTransformers  [^1] https://huggingface.co/sentence-transformers stands out as a state-of-the-art technique for generating sentence and text embeddings. Renowned for its ability to capture semantic similarities between documents, it is among the most popular choices for this task. There are many pre-trained models available for SentenceTransformers, all hosted on the Huggingface Model Hub [134]. The all-* models were trained on an extensive dataset of over one billion training pairs, making them suitable for general-purpose applications. In addition, there are Multi-QA models trained on 215 million question-answer pairs from diverse sources, including StackExchange, Yahoo Answers, and search queries from Google and Bing. Notable models in this category include multi-qa-MiniLM-L6-dot-v1, multi-qa-distilbert-dot-v1, and multi-qa-mpnet-base-dot-v1. The multi-qa-mpnet-base-dot-v1 model excels in semantic search performance, whereas the multi-qa-MiniLM-L6-dot-v1 model is optimized for speed. We chose to use the multi-qa-MiniLM-L6-dot-v1 model for our embedding needs because it is trained on question-answer data, which closely aligns with the nature of our dataset.
In BERTopic, there is a parameter nr_topics to control the number of topics, which merges topics after they have been created by supporting the creation of a fixed number of topics. However, it is recommended that the number of topics be controlled through a cluster model. We utilized the HDBSCAN [135] cluster model to perform this task. We improved the default representation of topics using the Countvectorizer  [^2] https://scikit-learn.org/stable/modules/generated/sklearn.feature_extraction.text.CountVectorizer.html. It helps to ignore infrequent words and increase the n-gram range. Following previous works [21, 17], we used the unigram and bigram models. Finally, we obtained 10 topics for SO data and 13 topics from the GitHub data. BERTopic includes a -1 topic to group documents that do not align well with the identified topics. This situation may arise when a document contains excessive noise, such as stopwords or irrelevant information, or does not strongly connect to the main themes identified by the model. The -1 topic serves as a ’catch-all’ for these outlier documents. However, since we preprocessed the data to eliminate such elements before applying BERTopic, our main topic began with the label -1. The scripts for data preprocessing and topic generation for SO and GitHub data are provided in the replication package [51]. We described our findings in the following Section.
4 Results
In this section, we present the analysis of the SWS-related data we obtained from SO and GitHub and the topics to answer our research questions.
4.1 RQ1: What topics are SWSs developers asking about on Stack Overflow, and which topics are raised in the GitHub projects?
Motivation: Scientific Workflow Systems (SWSs) development has unique characteristics that set it apart from traditional software development [59]. For example, SWSs developers need expertise in complex data management and the involvement of specialized software, high-performance resources, and hardware. SWSs should be adaptable to frequent changes, and adherence to community standards, reproducibility, and computational efficiency are crucial. Multidisciplinary team involvement and usage of domain specific languages are often necessary. These particularities make SWS development a complex endeavor, requiring a deep understanding of scientific domains and advanced software engineering practices, which are often unnecessary or overlooked in most conventional software development tasks. Hence, the challenges SWSs developers face will likely differ from those of traditional software development. Developers use Q&A websites to communicate both problems and solutions, and they rely on GitHub for version control and collaboration, including code review, project management, and collective software development. This research question aims to analyze invaluable data from SO and GitHub to identify the most common and pressing topics related to SWSs and to explore the relationships between these topics on both platforms. By pinpointing the challenges frequently encountered by the SWSs community, we can highlight which topics are gaining traction and which are proving challenging to address. Understanding these trends is crucial for supporting the SWSs community and guiding future development efforts.
Approach: We used BERTopic to identify the different topics developers discuss on SO and GitHub, as mentioned in Section 3. The first three authors (annotators) labeled the set of topics based on the overall theme of the posts. Each annotator individually inspected the top 15 keywords and a random sample of at least 30 posts from each topic to assign a title that best represents the content. The authors then discussed the labels for each of the 10 topics to reach a consensus on their titles. We followed a similar approach to label topics for the GitHub issues and pull requests. Additionally, we examined the most popular SWSs topics among developers. To investigate this, we used two complementary measurements of popularity that have been adopted in prior works [17, 109, 19, 136, 118]:
  1. 1.
    The average number of views (avg. views) of the post received from both registered and unregistered users serves as a valuable metric. A high view count implies significant interest and popularity among SWSs developers. This metric effectively measures community engagement by indicating how frequently a post is viewed.
  2. 2.
    The average score (avg. scores) of the posts received from the users serves as another important metric for measuring popularity. SO allows its members to upvote posts they find interesting and valuable and downvote posts to maintain high content standards. These votes are aggregated into a score, serving as a metric of the post’s perceived community value.
Table 2: The SWSs topics, keywords, and their popularity for SO data.
SL. Topic Keywords # Posts AvgView AvgSc
1 Workflow Creation and Scheduling dag, run, create, task, try, file, error, operator, use, docker. 10541 2656.19 1.95
2 Distributed Task Management task, worker, run, queue, start, work, app, celery worker, try, app. 9447 2616.23 3.04
3 Processor for Data Processing processor, use, file, flow, datum, json, flowfile, try, value, attribute. 4995 1281.33 0.79
4 Data Structures and Operations column, file, dask dataframe, datum, memory, try, code, array, function, read. 4703 1203.47 1.74
5 Workflow Execution workflow, job, action, hadoop, hive, run, spark, error, xml, file. 2020 1587.62 0.97
6 Managing Rules, Input, Output, and Files rule, file, output, input, wildcard, run, error, pipeline, sample, try. 1691 697.58 1.28
7 Task Dependencies and Management use, try, error, run, joint, pydrake, task, file, simulation, code. 1137 519.33 1.07
8 Data Transformation jolt, spec, output, input, value, transform, array, expect, use jolt, transformation. 448 513.69 0.57
9 Learning process, try, channel, file, output, script, pipeline, input, run, use. 381 729.80 1.12
10 Automation with Ansible ansible, role, collection, install, yml, module, instal, git, error, run. 256 3159.54 3.02
Results Obtained Using SO: Following the process described in sub-section 3.3, we identified 10 topics for SO data. Table 2 presents the obtained ten topic titles along with their associated keywords. It also shows the number of posts related to each topic and the popularity of these topics based on our metrics: views and scores received by developers on SO. As shown in Table 2, developers inquire about various topics in SWSs development, with the number of posts varying across topics. Below, we discuss these topics in greater detail.
1. Workflow Creation and Scheduling: This is the most dominant topic, representing 29.59% of posts. It involves defining and organizing a series of computational or data-processing tasks to achieve a particular scientific goal. Workflow creation entails specifying the individual tasks, their dependencies, and the required resources. Scheduling is the process of allocating these tasks to computational resources over time, ensuring efficient execution while considering factors like task priorities, resource availability, and deadlines.
One example of this topic with about 114k views is ’How to create a conditional task in Airflow’. In this case, the developer was trying to create a conditional task in an SWS based on a schema using ShortCircuitOperator and/or XCom to manage the condition but was unsure how to implement it and asked for a solution. This process is crucial for optimizing performance and resource utilization in complex scientific experiments and data analysis workflows. The prominence of this topic highlights the necessity of more intuitive ways to create complex workflows and manage task execution efficiently.
2. Distributed Task Management: This is the second most prominent topic, accounting for 26.52% of the posts. This topic discusses issues such as orchestrating complex task sequences, identifying pending tasks while ensuring fault tolerance and scalability, managing resources efficiently, optimizing performance, and handling memory allocation across distributed environments with large datasets. One notable example with 272K views is ’How to retrieve a list of tasks in a queue in Celery?’, where a developer executing a series of tasks sought help in retrieving a list of unprocessed tasks. Another example with 61K views is ’How do I use the @shared_task decorator for class-based tasks?’, where the developer needed guidance on decorating a class-based task. The high volume of posts on this topic indicates a need for improved task management to enhance the utilization of SWSs.
3. Processor for Data Processing: SWSs provide a rich set of processors covering a wide range of data integration, transformation, routing, and interaction tasks. Practitioners can design sophisticated data pipelines to meet their specific data processing needs by selecting and configuring the appropriate processors. This topic covers 14.02% of the posts. One example of this topic with about 27K views is ’How to use NiFi ExecuteScript processor with Python?’. ExecuteScript processor allows users to write and execute custom scripts within a data flow. This processor provides flexibility for handling complex processing logic that the built-in processors do not cover. Here, the developer tried to use the ExecuteScript processor with a simple Python script but failed to see the output and asked for help to explain how to use it with an example. Another example of this topic with 12K views is ’How to specify output filename with PutFile processor?’. Here, the problem was that the PutFile processor was saving files using a timestamp, but the expectation was to save them with the original filename. Thus, the developer is seeking a solution. This topic indicates developers’ challenges in understanding data processing processors.
4. Data Structures and Operations: This topic encompasses 13.20% of posts. In SWSs, data structures (i.e., arrays, dataframes, functions, and matrices) and operations (i.e., data ingestion, cleaning, transformation, analysis, visualization, storage, and parallel processing) are crucial for organizing and managing data. These components collectively facilitate the efficient processing and analysis of scientific data to derive meaningful insights. An example of this topic with 175K views is ’Make Pandas DataFrame apply() use all cores?’. The concern was that DataFrame.apply() is limited to a single core, causing a multi-core machine to underutilize its computing power. Thus, the developer was inquiring how s/he can use all the cores of a machine to run DataFrame.apply() on a dataframe to run in parallel. Another example with 47.5K views is ’Convert string to dict, then access key: values?. How to access data in a $<$class ’dict’$>$ for Python?’. Here, the developer faced issues accessing data from a dictionary and seeking a solution.
5. Workflow Execution: Workflow execution in SWSs entails orchestrating a predefined sequence of computational tasks to analyze scientific data, manage dependencies, allocate computational resources, and monitor progress. While workflow execution stands as a pivotal component within every SWS, developers encounter challenges in its implementation. 5.67% of posts within our dataset highlight this specific issue. An illustrative example in this topic, garnering about 42K views, is titled ’Stop Oozie workflow execution’. In this scenario, a developer initiated two jobs within a workflow but subsequently decided to terminate them. Despite halting the jobs, the workflow persisted. Consequently, the developer sought a command-line solution to cease the workflow execution effectively.
6. Managing Rules, Input, Output, and Files: In SWS, a rule defines a unit of work within a workflow, detailing how to generate output files from input files. Each rule typically includes directives for input, output, parameters, resources, and conditions, as well as the shell commands or scripts to be executed. This topic covered 4.75% of posts in our dataset and asked various issues related to defining and running rules, recommended practices for executing shell commands and scripts, and file naming using wildcards. For instance, a question with approximately 17K views is titled ’How to run only one rule in Snakemake?’. In this scenario, the developer created a workflow consisting of three rules: A, B, and C, where the output of rule A serves as the input for rule B, and the output of rule B serves as the input for rule C. Initially, the workflow ran smoothly, but the developer later wanted to run only rule C and encountered difficulties, thus sought help on SO.
7. Task Dependencies and Management: This topic explores issues such as managing complex inter-task dependencies, scalability, efficient resource allocation, intermediate data reuse, and concurrency. It also discusses challenges like implementing robust error handling and recovery mechanisms to maintain data integrity and ensure reproducibility through consistent environment management and version control. An example of this topic with around 12K views is ’How to reset Luigi task status?’, where a developer faced issues with a dependency chain ($a->b->c->d$), where task d was executed first, and task a was executed last. Then, the developer needed to rerun the same queue, changing the data. When rerunning the same queue, the Luigi SWS used historical data, so the developer sought a way to force the tasks to run without relying on this previous data.
8. Data Transformation: In SWS, this topic discusses converting data from one format or structure to another, enabling easier analysis, visualization, or integration with other datasets. This involves altering the shape or appearance of data objects while preserving their content and, for instance, using libraries like Jolt to transform JSON data from one format to another. This topic covers around 1.26% of posts of our SO dataset. One example of this topic, with around 10K views, is titled ’Is it possible to concatenate the values of JSON attributes using JOLT?’. In this instance, the developer inquired whether JOLT could be used to concatenate the values of three JSON attributes into a new attribute and also asked potential alternative approaches.
9. Learning: Questions in this topic are about how to do a particular task in SWS. This topic covers 1.07% of our posts. One example of this topic with around 6.4K views is ’How to run Python Script in Nextflow?’. The developer attempted to run an executable Python script in an SWS (NextFlow) but encountered an error stating No such file or directory, though it has the file. Subsequently, the developer asked for a solution for this issue on SO. Another example of this topic, with around 5K views, is titled ’How to generate the output directory for a Nextflow process within the Docker container?’. In this case, the developer ran two processes using Nextflow inside a Docker container. The first process successfully generated some plots, which were intended as input for the second process. However, the second process could not access the plots, prompting the developer to seek a solution in SO to make these plots available for the subsequent process. This topic highlights a significant gap in available documentation, tutorials, and other guidance resources for using SWSs.
10. Automation With Ansible: This is the least dominant topic, accounting for just 0.72% of posts. It focuses on the use of Ansible in SWSs. Ansible enhances SWSs by providing robust automation for environment setup, configuration management, workflow execution, and orchestration, ensuring scientific research consistency, reproducibility, and efficiency. One of the most viewed questions in this topic, with 111k views, is ’How to automatically install Ansible Galaxy roles?’ where Ansible is an open-source automation tool for deploying and configuring large-scale infrastructures [137]. In this question, the developer faced the challenge of manually downloading Galaxy roles on each machine and sought a method to automatically identify and download these roles when running Ansible on a new machine. We observed that many questions within this topic pertain to Ansible installation and role management. This trend likely stems from a lack of proper documentation or tutorials on using the tool effectively.
SWSs developers inquire about various aspects of SWSs development, including learning resources, automation, workflow creation, scheduling, data transformation, and task management. The most popular topic is distributed task management. This is because it deals with the complexities of coordinating and executing tasks across multiple systems, which is crucial for handling the large-scale data and computational workloads typical in scientific workflows.
Results Obtained Using GitHub: We identified 13 key topics developers encounter when working with GitHub data associated with SWSs following the process described in sub-section 3.3. Our analysis of GitHub issues and pull requests revealed that these challenges often stem from the inherent complexity and interdisciplinary nature of these systems. Table 3 presents a summary of the identified topics, their associated keywords, and a short description of each topic. Below, we provide a detailed description of each topic.
Table 3: Topics derived from SWSs on GitHub
SL. Topic Keywords Description
1 Errors and Bugs Fixing error, run, fix, bugs, case, tool change, add, file, test, task. Addressing errors and bugs in code, including fixes, version updates, and test cases.
2 Documentation documentation, change, add, test, version, license, code, task, file, case. Modifying and improving project documentation, including version updates and test cases.
3 Dependencies dependency, change, file, license, updated, build, tracking, branch, jdk. Managing and updating project dependencies and related files.
4 Data Structures and Operations array, dataframe, column, file, version, python, example, panda, conda, parquet. Working with data structures, such as arrays and dataframes, and related operations.
5 Automating Code Quality Checks precommitci, automatic update, precommitci start, end, rebase freeze, biotools. Implementing automated code quality checks and continuous integration processes.
6 Managing Releases release, created, generated, beep, documentation, bug fixes, chore-release, features. Handling the creation and management of software releases.
7 User Interface ui-ux, enhancement, slide, src, tutorial, visualisation, improvement, architecture. Enhancing the user interface and user experience of software projects.
8 Improving Code Quality flake, xref, fix, python, flake-import-order, error, add, remove, update, style. Improving code quality through refactoring, and error fixes.
9 Data Management and Task Scheduling task, scheduler, shuffle, data, partition, worker, disk, added, algorithm. Managing data and scheduling tasks within the workflow system.
10 Image Optimization optimisation, calibre, image actions, reduced, compression, filename, improvement, auto. Optimizing images for performance and efficiency.
11 Annotations annotation, type, related, add type, result, validated, fix, task, refactor, update. Adding and managing code annotations and type validations.
12 Kubernetes Deployments kubernetes, cluster, node, helm, install, chart, close, reproduce, reported, free, reopen, explain. Deploying and managing applications on Kubernetes clusters.
13 Maintaining Changelog changelog, update, add, release, fix, change, version, doc, file, missing, support. Maintaining and updating the project changelog with relevant changes and updates.
1. Errors and Bugs Fixing: This topic encompasses 46.2% of all issues and pull requests in the dataset, highlighting its significance. Addressing bugs and errors is crucial for ensuring the reliability and performance of SWSs. Bugs can disrupt functionality, cause unexpected behaviors, or undermine the user experience. For example, one issue involves ’fixing invocation serialization when no state is set’. Another such issue is ’fix bug for reduction of zero dimensional array’. Developers strive to create robust and reliable tools by prioritizing and resolving this topic, minimizing disruptions, and enhancing overall stability.
2. Documentation: This topic addresses a range of issues and pull requests crucial for maintaining and enhancing SWSs. It involves removing outdated content, offering vital support to users and contributors, and enriching documentation with practical examples. Additionally, it covers configuring tools within the workflow system, updating tutorials to reflect the latest features, setting up the necessary environments, and sharing best practices. For instance, in the case of setting up multiple Airflow instances under the same domain with different sessions, a developer faced challenges in installing and using multiple instances of an SWS and sought documentation for assistance by raising an issue ’Set up multiple Airflow instances under the same domain with different sessions’. Proper documentation provides essential guidance for both users and developers, facilitating a better understanding and implementation of workflows. Regular updates and clear communication about system changes help maintain a collaborative and productive environment, ensuring that all stakeholders are well-informed and equipped to work effectively within the system.
3. Dependencies: This topic gathers issues and pull requests related to the integration of libraries or other software dependencies in SWSs. One example is the issue ’Update elastic-transport requirement from $\leq 8.10.0$ to $\leq 8.11.0$, where elastic-transport, a Python library, serves as a transport layer and provides the infrastructure for HTTP communication, connection management, and request/response processing. SWSs utilize thousands of libraries like elastic-transport, pymongo, mypy, and conda, which can present various challenges during their implementation. Managing these dependencies ensures compatibility and smooth operation, preventing conflicts that can arise from version mismatches or outdated libraries. Effective dependency management, including clear documentation and automated installation scripts, is essential for maintaining a reliable and reproducible workflow environment.
4. Data Structures and Operations: This topic delves into the challenges and intricacies of data structures and operations, which are vital for managing complex datasets and orchestrating computational tasks in SWSs. These systems, such as Galaxy and Apache Airflow, exemplify the importance of handling diverse data types and datasets. Galaxy, for example, uses a variety of data structures, including arrays, dataframes, fastq, fasta, and files, to process biological data, emphasizing the need for managing and ensuring the compatibility of different data formats. Similarly, Apache Airflow leverages data structures and operations to facilitate data pipeline orchestration, as evidenced by issues like the error message ’NoneType’ object has no attribute ’replace’ while killing rabbit group task and the feature request ’Support setting inlets/outlets during task execution’, where a developer seeks to set datasets during task execution. To tackle this topic, developers strive to manage datasets efficiently, ensuring seamless data transformation, validation, and integration processes, which are critical for the effective operation of SWSs.
5. Automating Code Quality Checks: In SWSs, maintaining high code quality and consistency is essential for reliable data analysis and reproducibility. Tools like pre-commit ci are invaluable in this process, as they automate code quality checks and enforce coding standards. By running a series of predefined checks—such as linting, formatting, and security scans—automatically whenever code changes are made, pre-commit ci ensures that any issues are identified and addressed promptly. This mechanism, often called precommitci autoupdate, keeps the codebase aligned with the latest best practices and standards by automatically updating the checks. An example of this is seen in the pull request titled ’TS config and linting tweaks’, where pre-commit ci was used to enforce code quality standards. This pull request illustrates how automated checks can detect potential issues early in the development process, ensuring that the code adheres to the project’s guidelines before being merged. This pull request demonstrates how automated checks can catch issues early, ensuring the code adheres to the project’s guidelines before merging.
6. Managing Releases: Managing release is a crucial milestone that brings together various updates and improvements. When a new version is created, it signifies that a significant set of changes has been finalized and is ready for deployment. A generated release might follow, which involves automating the packaging and distribution of the updated software. Key elements of this process include addressing bug fixes to resolve issues reported by users and integrating features that enhance the software’s functionality. Additionally, a documentation release is often part of the process, ensuring users have up-to-date guides and references. Routine updates, or chore releases, might also be included to handle maintenance tasks such as dependency updates or configuration tweaks. For instance, ’Add note on fastapi to admin release notes’ exemplifies how significant updates and bug fixes are integrated into the software, reflecting the importance of a well-managed release process. Similarly, the issue ’Release Celery v5.3.5’ highlights how tracking and resolving bugs contribute to the overall quality and stability of the release.
7. User Interface: In SWSs, UI/UX design ensures that complex processes are accessible and manageable for researchers and scientists. A well-designed user interface (UI) and user experience (UX) can significantly enhance the usability of these systems, enabling users to seamlessly create, execute, and monitor workflows. Effective UI/UX in SWSs often involves intuitive layout and navigation, clear visualization of data and process steps, and interactive elements that facilitate user engagement and manipulation. By prioritizing user-centric design, these systems can reduce cognitive load, minimize errors, and improve the overall efficiency of scientific research. For example, features like drag-and-drop functionality for designing workflows or interactive dashboards for real-time monitoring can simplify intricate tasks and streamline workflows. Specific improvements such as ’Fix collection type drag and drop validation’ and ’Workflow invocation view improvements’ illustrate how targeted UI/UX enhancements can accelerate scientific discovery and innovation.
8. Improving Code Quality: In SWSs, maintaining high code quality is crucial for reliable and accurate computations. A common issue, known as the flake error, often arises from inconsistencies in code style or improper import order, as detected by tools like flake-import-order. For example, consider the issue ’Python package import collision’ in the Galaxy SWS. To resolve these problems, developers need to fix flake errors by adhering to consistent coding practices and arranging imports correctly. The xref mechanism is also vital, enabling the addition (xref add) and updating (xref update) of cross-references across modules and workflows, which helps track dependencies and maintain a clear structure. When an xref fix is necessary, correcting the error and updating the cross-references is important. Additionally, developers may need to remove outdated or conflicting components and pin specific library versions to ensure consistency and reproducibility. By implementing these practices, SWSs can become more reliable and maintainable, reducing errors and ensuring efficient execution of scientific computations.
9. Data Management and Task Scheduling: Effective data management and task scheduling are pivotal for optimizing performance and resource utilization in SWSs. The core concept revolves around breaking down complex tasks into smaller, manageable units, which are then assigned to workers by a scheduler. This scheduler uses sophisticated algorithms to determine the optimal sequence and allocation of tasks, ensuring efficient execution. One critical aspect is the partitioning of data, which involves dividing large datasets into smaller chunks or partitions. These partitions are distributed across workers, facilitating parallel processing and reducing the computational load on individual nodes. During the workflow, tasks may require data shuffling, where data is moved between workers based on specific criteria, ensuring that each worker has the necessary data to complete its assigned tasks. Data may be passed through the system’s memory or written to disk, depending on the workflow’s requirements and the system’s architecture. This topic discusses various aspects of effective scheduling and data management. For instance, one of the practitioners of SWSs raised an issue ’New jobs are not getting scheduled due to the error ”InternalError galaxy.jobs.handler in __handle_waiting_jobs’. Another pull request raised by other practitioners is ’[Dask.order] Ignore data tasks when ordering’.
10. Image Optimization: In SWSs, image optimization enhances performance and storage efficiency. Optimizing images involves reducing their size through compression techniques without significantly compromising quality. This process, known as image compression, can reduce the file size, making it easier to handle large datasets and speeding up data transfer and processing times. Image optimization techniques can also identify the more important image components and discard less important components. Additionally, improving filenames with descriptive and standardized formats ensures better organization and accessibility within the system. Tools like Calibre can be utilized for automated image actions, such as resizing and format conversion, further streamlining the workflow. By implementing these optimization strategies, scientific workflow systems can operate more efficiently, facilitating quicker data analysis and more effective resource management. One example of this topic is ’Auto Compress Images by Calibre’s image-actions’.
11. Annotation: In SWSs, annotation involves the addition of metadata or comments to various workflow components, such as data, processes, or results. These annotations can take the form of descriptions, tags, explanations, or contextual information, which collectively enhance the clarity, reproducibility, and usability of scientific workflows. This topic explores annotation issues, such as the problem where ’annotations somehow appear in workflow jobs’. In certain SWSs, like Celery, annotations play a significant role in adding extra information to tasks, thereby influencing their behavior or providing additional context. For example, the ‘max_retries‘ annotation in Celery specifies the maximum number of times a task should be retried if it fails. A specific issue related to this is ’Add annotations to minimize differences with celery-aio-pool’s tracer.py’
12. Kubernetes Deployments: In SWSs, efficient deployment and management of applications across Kubernetes clusters are essential. Helm, a powerful package manager for Kubernetes, plays a vital role in this process by using Helm charts to define, install, and upgrade complex applications. When setting up a new cluster, it’s essential to ensure that all nodes are configured correctly and Helm charts are deployed as intended. This topic discusses issues such as how to reproduce an issue reported by users and how to track and manage these effectively, from filing the initial report to potentially reopening it if the problem persists and eventually closing it once resolved. One such issue is ’Make helm chart compatible with $apache/nifi\geq 1.19.0$in Kubernetes cluster’.
13. Maintaining Changelog: Maintaining an accurate changelog is essential for any software project, including SWS. A detailed changelog provides a comprehensive record of all changes made in each release, helping both users and developers stay informed about new features, bug fixes, updates, or deprecations. It also includes documentation updates that clarify usage or functionality. Without a proper changelog, users might overlook important updates or changes, leading to confusion or issues during implementation. Therefore, consistently adding entries to the changelog for every release is a best practice, ensuring that all modifications are clearly documented. This topic addresses various issues related to changelogs, such as the problem of a missing changelog entry in an SWS project, exemplified by the issue ’Add missing changelog’.
Tables 2 and 3 present topics related to SWSs but from different sources—SOdata and GitHub. A notable similarity between the two tables is the presence of topics focused on data structures and operations, as well as task management and scheduling. The ’Data Structures and Operations’ topic in both tables highlights the significance of handling dataframes, arrays, and related data manipulation tasks within SWS. Additionally, task management is a recurring theme, with the ’Workflow Creation and Scheduling’ from SO data and ’Data Management and Task Scheduling’ from GitHub, both emphasizing the orchestration and execution of tasks within workflows. Moreover, both tables reference topics related to errors and bugs, with GitHub specifically addressing ’Errors and Bugs Fixing’, while SO topics such as ’Workflow Execution’ and ’Managing Rules, Input, Output, and Files’ implicitly involve error handling and troubleshooting. This overlap underscores the importance of managing data, tasks, and errors in SWSs across different platforms.
4.2 RQ2: What types of SWSs development questions are asked on technical forums?
Motivation: Building upon identifying key topics in RQ1, a critical question emerges: What type of questions are SWSs developers asking? Answering this is crucial as it deepens our understanding of the role SO plays for SWSs developers. This investigation shifts the focus from just what developers ask to why they ask it. By analyzing the types of posts created within each SWS topic, we aim to uncover the nature of the challenges developers face during SWSs development. Prior research [16, 17] highlights that developers typically pose different types of questions—such as ’How’, ’Why’, and ’What’—to address specific challenges. This detailed analysis will not only highlight the common problems encountered but also expose the underlying complexities and pain points within the SWSs development process. Categorizing these questions will provide insight into developers’ cognitive processes and problem-solving strategies, helping to identify areas where additional support or resources are needed. Ultimately, this work offers a deeper understanding of the reasons SWSs developers turn to SO instead of relying solely on official documentation.
Approach: To achieve that, we followed a similar approach used by prior works to identify the types of posts on SO [16, 17, 18]. In particular, we randomly sampled SO posts from each of the ten topics with a confidence level of 95% and a confidence interval of 5%. Our random sample size of each topic yields a total of 2933 posts: 371 Workflow Creation and Scheduling, 370 Distributed Task Management, 357 Processor for Data Processing, 356 Data Structures and Operations, 323 Workflow Execution, 314 Managing Rules, Input, Output, and Files, 288 Task Dependencies and Management, 208 Data Transformation, 192 Learning, and 154 Automation with Ansible. We reviewed each question’s title and body to determine its type (at least two authors review each post). We manually analyzed and categorized each sample question. Each question is assigned a single label; if it could fall under multiple labels, we selected the most relevant one. Finally, each annotator labeled the post using one of the following types that were used by prior works [16, 17]. In cases where annotators couldn’t reach a consensus, they revisited and discussed the questions together to achieve an agreement. Despite the annotators’ excellent comprehensive understanding of SO data, we carefully review all the labeled data from the replication package of [17] where a similar task was done for chatbot development, together to ensure a shared understanding of the labeling process. Overall, the annotators achieved almost perfect agreement [138], with a kappa value of 0.82, on the 2,933 classified posts. This high kappa value signifies a very strong level of consistency among the annotators, indicating minimal discrepancy in the assessments. Such a result underscores the reliability of the classification process and suggests that the annotated data is robust, with the raters largely aligning in the judgments.
  • How: This category includes posts where developers seek guidance on specific methods or techniques to accomplish a particular task [16, 17]. Unlike ’Why’ posts that inquire about underlying reasons or explanations, these posts are goal-oriented. Developers ask for precise, step-by-step instructions to achieve their specific objectives (e.g., ’How to remove default example DAGs in Airflow’).
  • Why: This type includes posts where developers inquire about certain phenomena’ reasons, causes, or purposes [16, 17]. ’Why’ posts often pertain to troubleshooting, where developers seek explanations for specific behaviors or issues (e.g., ’Why does Celery need a message broker?’).
  • What: This category includes posts where developers seek specific information [16, 17]. These inquiries are often aimed at clarifying doubts to make more informed decisions. For example, a developer asked, ’What is the best approach for adding cron jobs (scheduled tasks) for a particular service in Docker Compose?’ in SO.
  • Others: We assigned this type to posts that do not fit into any of the above-described types (e.g., ’Understand the Notify and Wait process in NIFI in my flow?’).
To evaluate the quality of our classification of the random sample, we utilized Cohen’s Kappa [139] to quantify the level of agreement among the annotators. This statistical measure helps us assess the consistency and reliability of our categorization process.
Table 4: SWSs posts types on Stack Overflow
Topic % How % Why % What % Other
Workflow Creation and Scheduling 51.75 21.29 23.99 2.96
Distributed Task Management 61.89 19.46 17.84 0.81
Processor for Data Processing 63.31 13.45 18.77 4.48
Data Structures and Operations 58.71 13.76 24.16 3.37
Workflow Execution 44.89 21.98 28.79 4.34
Managing Rules, Input, Output, and Files 65.92 16.24 16.56 1.28
Task Dependencies and Management 60.42 20.49 17.71 1.38
Data Transformation 76.92 10.58 9.13 3.37
Learning 71.35 14.58 14.06 0.0
Automation with Ansible 54.55 20.13 24.68 0.65
SWSs (all) 60.97 17.2 19.57 2.26
Results: Table 4 shows the percentage of the posts types for each SWSs topic.
Figure 6: Types Distribution of SO Posts
The analysis of question distributions across various topics in SWSs reveals distinct patterns that underscore developers’ needs and knowledge gaps. The predominant focus on ’How’ questions, averaging 60.97% across all topics, indicates a significant demand for procedural guidance among users. For instance, topics such as ’Workflow Creation and Scheduling’ (51.75%) and ’Distributed Task Management’ (61.89%) demonstrate users’ strong interest in understanding the execution of specific tasks. Similarly, ’Processor for Data Processing’ (63.31%) and ’Data Structures and Operations’ (58.71%) highlight the necessity for detailed instructions and methodologies. Conversely, the ’Why’ questions, constituting an average of 17.2%, reflect users’ curiosity about the underlying reasons behind certain actions, with ’Workflow Execution’ and ’Task Dependencies and Management’ topics showing notable percentages (21.98% and 20.49%, respectively). ’What’ questions, accounting for 19.57% on average, indicate users’ need for clarification on the specifics of tasks or concepts, as seen in ’Workflow Execution’ (28.79%) and ’Automation with Ansible’ (24.68%). The ’Other’ category, with a minimal average of 2.26%, suggests that most user inquiries are well-categorized within the primary question types. Figure 6 shows the summary of types distribution for our analysis.
We observed that the ’How’ type consistently dominates across various domains. For instance, ’How’ questions constitute 61.8% of chatbot-related posts and 59% of mobile development posts, closely aligning with our domain’s 60.97%. The second most common question type in chatbot and mobile development topics is ’Why’ (25.4% and 29%, respectively). Interestingly, the second most common question type in our domain is ’What’, whereas ’Why’ questions rank third. Conversely, ’What’ questions are the third most frequent in the chatbot and mobile development domains.
The higher percentage of ’What’ type questions compared to the ’Why’ type in SWSs posts on SO can be attributed to the practical and task-oriented nature of scientific computing. Users in this domain often seek clarity on specific functionalities, features, or tools necessary for their workflow, which drives a higher frequency of ’What’ inquiries. These questions typically focus on understanding particular components, commands, or processes within their workflows. In contrast, ’Why’ questions, which delve into the rationale or underlying principles, are less frequent because users are more immediately concerned with the practical implementation and troubleshooting of their tasks. This trend reflects the urgency to solve concrete problems and efficiently utilize tools and methods in SWSs.
4.3 RQ3: Which topics are most difficult to answer?
Motivation: Given our understanding of different topics obtained from SO and GitHub, our next step is to investigate the difficulty of answering posts, resolving issues, and pull requests within each identified topic. By determining whether some topics are more challenging to address than others, we can identify areas that require increased community support. Additionally, this analysis can help us highlight topics where there may be a need for improved tools or frameworks to assist developers in tackling SWSs development challenges.
Approach: We assessed the difficulty of each topic using two distinct metrics. For SO, we considered the percentage of posts within a topic that lack accepted answers and the median time, in hours, for an answer to be accepted. On GitHub, our metrics include the percentage of unresolved issues and pull requests, along with the median time taken to resolve them. This approach is consistent with methods used in previous studies [16, 19, 20, 21, 22, 17]. The metrics are described in detail below:
Metrics for SO:
  1. 1.
    The percentage of posts of a topic without accepted answers (w/o accepted answers). For each SWSs topic, we measured the percentage of posts without accepted answers. Although a post can receive multiple answers, only its author can mark one as accepted if it satisfactorily resolves the original question. Consequently, topics with fewer accepted answers are considered more challenging [17, 16, 19].
  2. 2.
    The median time for an answer to be accepted (median time to answer (hrs.)). We measured the median time, in hours, for posts to receive an accepted answer. This metric is based on the time the accepted answer is created, not when it is marked as accepted. A longer time to receive an accepted answer indicates that the post is more challenging [17, 16, 19].
Metrics for GitHub:
  1. 1.
    The percentage of unresolved issues and pull requests of a topic (w/o solutions). The unresolved issues and pull requests percentage within a specific topic indicates the proportion of issues and pull requests that remain unresolved within that particular topic area. For each SWSs topic, we measured the percentage of issues and pull requests that are not closed (when a developer closes an issue or pull request in a software development context, it typically means that they have addressed the problem or implemented the requested feature and believe the issue or pull request is resolved). Therefore, topics with open state are considered more difficult.
  2. 2.
    The median time for an issue or pull request to be resolved (Median Time to Resolve(hrs.)). We calculate the median resolution time in hours for issues and pull requests, considering the duration between the issue’s or pull requests’ creation and closing times. This metric focuses explicitly on closed issues or pull requests since open issues or pull requests lack a closing time. The greater the duration it takes to be resolved, the more challenging it is considered.
Table 5: The difficulty per topic on SO data
Topic Posts w/o Accepted (%) Median Time (h)
Workflow Creation and Scheduling 63.32 9.09
Distributed Task Management 61.22 9.32
Processor for Data Processing 59.53 4.77
Data Structures and Operations 57.71 6.88
Workflow Execution 68.50 23.58
Managing Rules, Input, Output, and Files 52.43 7.73
Task Dependencies and Management 47.13 7.06
Data Transformation 34.08 1.81
Learning 44.74 6.82
Automation with Ansible 58.98 4.49
Results: Table 5 presents the percentage of questions without accepted answers and the median time (in hours) to receive an accepted answer for each identified topic of SO data. The topic of Workflow Execution has the highest percentage of unanswered questions and also requires the longest median time to receive an answer. To understand the high rate of posts without accepted answers (68.50%), we examined the posts within this topic. We discovered that posts without accepted answers often received low scores (666 posts have a score of zero) from developers on Stack Overflow. This may be due to unclear or poorly formulated questions, reducing the likelihood of receiving an accepted answer. Additionally, we explored why this topic has a significantly higher median response time. We found that questions on this topic often involve complex tasks such as chaining jobs, managing processes, running specific jobs on particular nodes, and integrating tools in SWS. These tasks require rigorous testing and workflow implementation as well as execution in SWS, which significantly increases the time needed to provide an accepted answer.
We also found that Workflow Creation and Scheduling and Distributed Task Management topics have low acceptance rates and high median time requirements to achieve an accepted answer. By exploring the posts, we identified 3,078 and 2,228 posts with zero or negative scores, indicating low-quality questions. The high median time for answers in these areas is due to the highly specific nature of the questions, which are often tied to individual setups, tools, and requirements. Additionally, reproducing these issues without a similar environment is difficult, complicating the process of diagnosing and providing solutions. Many times, offering a meaningful answer demands a thorough understanding of the specific infrastructure and software versions in use.
Conversely, posts related to Data Transformation boast the highest percentage of accepted answers and a remarkably low median time of just 1.81 hours to receive an accepted answer. These posts often focus on how to perform data transformations using specific technologies (e.g., How to convert string to Datetime with specific format using jolt transformation?) or transforming JSON, array, or vector data. This topic’s close connection to traditional software development likely explains why SO respondents tend to answer these questions more quickly and frequently.
To gain a comprehensive understanding of SWSs posts, we aim to examine whether there is a statistically significant correlation between their difficulty and popularity. Specifically, we used the Spearman Rank Correlation Coefficient [140] to assess the relationships between two popularity metrics (average views and average scores) and two difficulty metrics (percentage without accepted answers and median time to answer in hours).
Table 6: Correlation of topics popularity and difficulty.
Correlation Coeff./p-value Avg. Views Avg. Score
% w/o Accepted Answers 0.673/0.033 0.458/0.183
Median Time to Answer(Hrs) 0.144/0.691 -0.053/0.884
We chose Spearman’s rank correlation because it makes no assumptions about the normality of the data distribution. As shown in Table 6, the analysis of correlations indicates a significant positive relationship between the percentage of unanswered questions and average views on SO, suggesting that topics with more unanswered questions tend to attract greater interest. This phenomenon can be attributed to developers frequently tackling intricate or specialized subjects that fewer people are knowledgeable about, which intrigues users facing similar challenges seeking solutions or insights. Moreover, unanswered questions often foster community engagement as users monitor them in anticipation of potential answers or opportunities to demonstrate expertise. For instance, the question How to retrieve a list of tasks in a Celery queue? has garnered over 273K views despite lacking an accepted answer. We found many questions with numerous views remain unanswered or lack accepted solutions. However, there is a very weak positive and statistically insignificant correlation between the median hours to answer and average views, suggesting no meaningful relationship. Similarly, there is a moderate positive but statistically insignificant correlation between the percentage of unanswered questions and average score, and a very weak negative and statistically insignificant correlation between the median hours to answer and average score. These results imply that while the percentage of unanswered questions may influence viewership, the time taken to answer and the average scores are not significantly correlated with either difficulty or popularity metrics.
Table 7: Unresolved Issues and Median Time for Different SWSs Topics on GitHub
Topic Name Unresolved (%) Median Time (h)
Errors and Bugs Fixing 9.62 45.68
Documentation 7.18 36.33
Dependencies 0.15 51.60
Data Structures and Operations 19.78 105.21
Automating Code Quality Checks 0.76 0.20
Managing Releases 0.62 4.70
User Interface 7.86 17.19
Improving Code Quality 4.38 8.43
Data Management and Task Scheduling 12.45 61.44
Image Optimization 0.47 0.91
Annotations 13.53 22.86
Kubernetes Deployments 27.27 84.46
Maintaining Changelog 5.34 4.44
Table 7 shows the percentage of unresolved issues and pull requests and the median time required to resolve them. The analysis reveals that Data Structures and Operations and Kubernetes Deployments are among the most complex topics within scientific workflow systems on GitHub. With 19.78% of issues and pull requests unresolved and a median resolution time of 105.21 hours, Data Structures and Operations pose significant challenges, likely due to the intricate nature of managing large datasets and performing complex operations. Similarly, Kubernetes Deployments, with the highest unresolved rate of 27.27% and a median resolution time of 84.46 hours, highlight the complexity of deploying and managing applications in distributed environments, which require specialized expertise and thorough troubleshooting.
Another challenging area is Annotations, where 13.53% of issues remain unresolved, and the median time to resolve them is 22.86 hours. The difficulties here stem from the precision needed in implementing type validations and annotations, which are crucial for ensuring code quality but are often complicated and time-consuming to address. Data Management and Task Scheduling also ranks high in difficulty, with 12.45% unresolved issues and a median resolution time of 61.44 hours, reflecting the challenges in orchestrating and managing tasks within large-scale workflows.
On the other end of the spectrum, topics such as Automating Code Quality Checks and Image Optimization are relatively easy to manage, with very low percentages of unresolved issues (0.76% and 0.47%) and quick resolution times (0.20 and 0.91 hours). These tasks are well-supported by automation tools, making them straightforward to address. Managing Releases and Maintaining Changelog also fall into the easier category, benefiting from structured processes and automation that contribute to low unresolved rates and swift resolutions.
The analysis of difficulty across various SWSs topics on SO and GitHub reveals distinct patterns. On SO, topics like Workflow Execution (68.50% unresolved, 23.58 hours median time) and Workflow Creation and Scheduling (63.32%, 9.09 hours) have higher unresolved post percentages, indicating their complexity. In contrast, Data Transformation (34.08%, 1.81 hours) seems easier for developers. Similarly, on GitHub, Data Structures and Operations (19.78% unresolved, 105.21 hours) and Kubernetes Deployments (27.27%, 84.46 hours) present more challenges, with prolonged issue resolution times. Meanwhile, topics like Automating Code Quality Checks (0.76%, 0.20 hours) are resolved swiftly. This analysis highlights the diverse difficulty levels developers encounter in SWS, depending on the topic.
5 Discussion & Implications of our Findings
In this section, we examined the evolution of SWSs topics and compared our results with those of previous studies. Subsequently, we analyzed the data to identify the dominant topics across various platforms and discuss the significance of our findings.
5.1 SWSs Topics Evolution
SWS is an emerging topic that is attracting increasing attention from developers across various domains, including software engineering [141], machine learning [142], bioinformatics [143], and data analysis[59]. To examine the evolution of SWSs, we present Figure 1, which shows the consistent growth of SWSs over time. The upward trend in the number of posts, issues, and pull requests indicates that SWSs topics are gaining significant attention from the community.
Figure 7: SWSs topics evolution over time on SO
To better understand the evolution of the different SWSs development activities, we measure the absolute growth of each of the ten topics over time shown in Figure 7. The data reveals a significant upward trend in the number of posts related to most topics, with ’Workflow Creation and Scheduling’ experiencing remarkably rapid growth, especially after 2015. This surge can be attributed to technological advancements that make workflow automation more accessible and essential for businesses. Tools and platforms like Apache Airflow, Kubernetes, and others have gained popularity, leading to more discussions. The proliferation of tools and platforms for workflow automation, coupled with a strong community focus on knowledge sharing, has further fueled this interest. Overall, the data underscores the critical importance of workflow and task management in enhancing productivity and efficiency in modern business environments.
Figure 8: SWSs topics evolution over time on GitHub
The absolute growth of GitHub 13 topics over time is shown in Figure 8. The evolution of topics related to SWSs reveals significant trends and shifts over the years. Initially, in 2009, the focus was primarily on documentation and error and bug fixing, with minimal attention to other areas. As the years progressed, there was a noticeable increase in the emphasis on ’Documentation’ and ’Errors and Bugs Fixing’, peaking around 2020. This indicates a growing recognition of the importance of thorough documentation and addressing errors in SWSs.
From 2015 onward, there was a substantial rise in topics related to ’Data Management and Task Scheduling’, ’Data Structures and Operations’, and ’Automating Code Quality Checks’, highlighting a growing focus on improving the efficiency and quality of SWSs. Notably, ’Data Structures and Operations’ saw a significant increase in attention, reflecting the importance of optimizing data handling within these systems. Similarly, the rise in ’Automating Code Quality Checks’ and ’Managing Releases’ demonstrates an evolving emphasis on maintaining and improving code quality and managing software releases. ’Image Optimization’ and ’Kubernetes Deployments’ emerged as relevant topics in later years, indicating the integration of new technologies and practices into SWSs. The data also show that ’User Interface’ and ’Improving Code Quality’ topics, while initially less prominent, have gained traction over time, underscoring the growing importance of user experience and code quality in SWSs.
Overall, the data reflects a broadening focus from basic documentation and error correction to more complex areas such as data management, code quality, and technological advancements, illustrating the evolving needs and priorities in the development of SWSs.
5.2 SWSs compared to Other SE Fields
SWS is an emerging field, making it essential to investigate how its topics compare to discussions in other software engineering domains such as chatbot [17], quantum software engineering [21], mobile apps [16], big data [19] and security [20]. To answer this question, we examined the difficulty and popularity of SWSs topics and compared them to other disciplines using data from similar studies on SO focused on chatbots, quantum software engineering, mobile apps, big data, and security. These domains have been previously studied and provide a useful benchmark. This comparison is crucial for identifying trends and gaps, allocating resources, and fostering innovation. Since these studies were conducted at different times, we used their reported keywords to construct an updated dataset and calculate the popularity and difficulty metrics for each field. This approach allows us to get insight into the current landscape and future directions of SWSs about other key areas in software engineering.
Table 8: Comparison of popularity and difficulty between different fields
Metrics SWSs Chatbot QSE WebApp Mobile Security BigData
# of Posts 35,619 7104 634 14075 2474527 169865 194356
Avg ViewCount 2119.46 996.46 719.06 2409.98 2792.04 3462.07 2799.27
Agv Score 1.96 0.78 1.10 2.12 2.40 2.73 1.77
Avg AnswerCount 1.09 1.06 1.08 1.02 1.45 1.41 0.84
% w/o Answers 60.03 67.69 60.57 64.31 53.21 52.33 58.83
Med. TimeToAns (Hrs) 7.67 16.16 9.17 11.38 0.91 1.75 3.26
Table 8 shows the result of the popularity and difficulty metrics among the seven fields. The metrics reveal several compelling reasons for increased focus on the development of SWSs. Despite having 35,619 posts and a high average view count of 2,119.46, indicating substantial interest, the community engagement for SWSs remains moderate. When compared to fields like Chatbot and QSE, SWSs perform better in terms of view count but still has a high percentage of unanswered posts (60.03%), similar to Chatbot (67.69%) and QSE (60.57%). The median time to answer for SWSs is 7.67 hours, which is slower than more mainstream technologies like Mobile (0.91 hours) and Security (1.75 hours) but faster than Chatbot (16.16 hours) and WebApp (11.38 hours). Additionally, SWSs have a lower average score (1.96) compared to fields like WebApp (2.12) and Security (2.73), indicating room for improvement in the quality of responses. Given the critical role that SWSs play in automating and optimizing complex scientific processes, enhancing community support, improving documentation and resources, and implementing measures to ensure quicker and higher quality responses are essential. This focused development can bridge the support gap, promote more effective utilization, and drive scientific research and innovation advancements.
5.3 Implications
The results of our study provide valuable insights that can help the SWSs community prioritize their efforts on the most critical challenges in SWSs development. Our analysis of SO and GitHub data reveals several key implications for both the development and utilization of these systems. Below, we outline how our findings can guide developers, researchers, tool developers, and educators in enhancing the practice and learning of SWSs development.
Implications for Practitioners: The analysis of SO and GitHub data offers valuable insights for practitioners working on SWSs. Here’s how these findings can guide their efforts:
  • Addressing Complex Challenges: Both Workflow Execution and Data Structures and Operations are particularly challenging, with high percentages of unresolved posts and long resolution times on both platforms. Practitioners should prioritize developing robust support mechanisms and optimizing data workflows to improve system performance.
  • Improving Workflow Management: The significant difficulty in Workflow Creation and Scheduling, as shown by unresolved posts and resolution times, suggests a need for better tools and processes. Enhancing these areas could lead to increased productivity and smoother operations.
  • Task Dependencies: Despite fewer unresolved issues, task dependencies remain a challenge. Practitioners should focus on simplifying task management to avoid complications in workflow execution.
  • Focusing on Deployment and Dependencies: With the highest unresolved issue rate and lengthy resolution times on GitHub, Kubernetes deployments is a significant pain point. Practitioners should seek specialized training or community support to overcome these challenges. Dependencies and Data Management also show prolonged issue resolution times and require careful planning and resource allocation to prevent delays in project timelines.
  • Enhancing Documentation and UI: Although documentation issues are less frequent, they take considerable time to resolve. Practitioners should invest in creating comprehensive documentation to assist users and reduce issue resolution times. Improving the user interface could make SWSs more accessible, especially for less technical users, reducing the number of related issues and enhancing the overall user experience.
In summary, practitioners should prioritize challenging areas like workflow execution, data management, and Kubernetes deployments, while also leveraging automation and enhancing documentation to improve the effectiveness and scalability of SWS.
Implications for Researchers: The analysis of SO and GitHub data provides several key insights that can guide researchers working with SWS. Notably, challenges in workflow execution, data management, and Kubernetes deployments stand out as significant pain points, requiring researchers to prioritize these areas to ensure efficient and scalable workflows. Leveraging automation tools like Ansible can greatly enhance efficiency, reducing manual effort and improving reproducibility. Additionally, the importance of robust documentation cannot be overstated, as it facilitates collaboration and knowledge transfer within the research community. By addressing these key areas, researchers can optimize their workflows and achieve more reliable and impactful research outcomes.
Implications for Educators: The analysis of SO and GitHub data and the distribution of post types on SO highlights several key areas where educators can enhance their teaching strategies for SWSs. The significant focus on ’how-to’ questions across topics like data processing and task management indicates a strong need for practical, hands-on learning experiences that address the operational challenges students will face. Additionally, the complexity and prolonged resolution times in areas like workflow execution, data management, and Kubernetes deployments suggest that these topics should be prioritized in the curriculum, with in-depth coverage and applied projects. Educators should also emphasize the importance of understanding the ’why’ behind processes, encouraging critical thinking alongside technical skills. By integrating these insights into their teaching, educators can better prepare students for real-world challenges in SWSs development and research, equipping them with both the practical know-how and the conceptual understanding needed for success.
Practitioners, developers, educators, and researchers can consider many aspects when deciding where to focus their efforts. Nevertheless, our study underscores the critical challenges and opportunities in developing and using SWSs. By addressing the identified issues and leveraging the insights from our analysis, stakeholders can enhance the effectiveness and impact of SWSs, ultimately advancing scientific research and innovation.
6 Threats to Validity
Threats to validity refer to factors or influences that can compromise the accuracy, reliability, or generalizability of study findings. Validity of research is concerned with the question of how the conclusions might be wrong, i.e., the relationship between conclusions and reality [144]. These threats can introduce errors, biases, or limitations that affect the validity and credibility of research results. Identifying and addressing these threats is essential to ensure the robustness and trustworthiness of research findings. In our study, we acknowledge the presence of the following potential threats.
External Validity: refers to the extent to which research findings can be generalized to broader populations or situations [145]. In this work, we analyzed Stack Overflow and GitHub repositories to understand the challenges of SWSs development. First, we used SO data to identify the challenges faced by developers of SWSs. However, this data might not encompass all the difficulties experienced by developers, as there could be subtle aspects rarely discussed. To address this limitation, we also examined prevalent issues and pull requests using GitHub, identifying common topics discussed there. Developers may also communicate their discussions through other media (e.g., mailing lists and SWSs forums). Future work considering other data sources may complement our study. Furthermore, another issue arises from the limited selection of only 11 SWSs based on their high prevalence on Stack Overflow. This selection may not represent the broader population of SWSs, as it focuses only on the most discussed systems, potentially overlooking less popular or emerging SWSs that could exhibit different characteristics or trends. The findings derived from these 11 systems may not generalize to all SWSs, especially those that are less documented or used by smaller, specialized communities. Consequently, the conclusions drawn might reflect biases inherent in the selection process, limiting the applicability of the results to the broader landscape of scientific workflow systems.
Internal Validity: pertains to the causal relationship between a treatment and its outcome [145]. In this study, we utilized topic models (BERTopic) to group related SO posts, and GitHub issues and pull requests, assuming that similar textual content would be present within the same clusters. However, alternative methods could result in different clusters of posts and issue reports. To ensure the quality of the clusters, we manually reviewed the generated topics, merging similar ones if necessary and assigning meaningful labels to each cluster. Another potential concern is selecting the optimal number of topics. In BERTopic, though there is a parameter nr_topics to control the number of topics, we controlled the number of topics through a cluster model. We used the HBDSCAN cluster model to perform this task. We improved the default representation of data using the Countvectorizer. It helps to ignore infrequent words and increase the n-gram range.
The subjectivity involved in labeling posts types poses a potential threat to the validity of our results. To mitigate this, we conducted three independent classifications and assessed the interrater agreement using the Cohen-Kappa test, which demonstrated a strong level of consistency among the annotators.
Construct Validity: involves the relationship between theory and observation [145]. A potential concern is the accuracy of labeling automatically generated topics, as the assigned names may not accurately capture the essence of the related posts, issues, and pull requests. To mitigate this threat, three authors independently reviewed the keywords and examined over 30 randomly selected posts/issues/pull requests for each topic. We then engaged in discussions to reach a consensus on a label that best represents the content of each topic.
Additionally, we evaluated the difficulty of SWSs topics by considering the percentage of unresolved posts, issues, and pull requests, as well as the median time to resolution. These metrics, while useful, may pose a threat to construct validity. Other potential indicators, such as the median time to the first answer, the number of answers per post (for Stack Overflow), and the number of comments, issue labels, review comments, and review cycles (for GitHub), could also be explored. However, the metrics we used are consistent with those applied in similar studies [17, 109, 136, 118, 16, 20].
7 Conclusion
This paper explores the challenges faced by developers of scientific workflow systems by analyzing relevant Stack Overflow posts and examining GitHub issues and pull requests. The findings reveal that developers encounter traditional software engineering challenges, such as data structures, operations, and dependency management, as well as unique challenges specific to SWSs, including workflow creation and task scheduling. Some SWS-related areas (e.g., workflow creation, execution, task scheduling, data structures, and operations) are gaining the highest attention from developers while being very challenging to them. We believe our findings provide valuable insights for the SWSs community by identifying areas that require focused research, particularly in addressing the most challenging aspects of SWSs development. Additionally, our results can assist SWS platform owners in gaining a better understanding of the challenges developers face and implementing strategies to address these issues effectively. Educators in the SWSs domain can also benefit from this study by incorporating the identified topics and their associated difficulties into their course design, ensuring that their courses are aligned with the real-world challenges encountered by developers.
Our study provides valuable insights for researchers and practitioners in the SWSs community, shedding light on the challenges faced by SWSs developers. However, we plan to extend our research by examining developer discussions across SWSs forums to draw more accurate and generalizable conclusions. Our future work will involve a comprehensive investigation of all SWSs forums, repositories, and posts to gain deeper insights into the various issues developers encounter and their strategies for resolving them.
Acknowledgments
This research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), and by the industry-stream NSERC CREATE in Software Analytics Research (SOAR).
Data Availability Statements
Our replication package can be found in our online appendix [51].
Conflict of Interest
We have no conflict of interest.